# Robust Radar Mounting Angle Estimation in Operational Driving Conditions

### v1

@@ Abstract
The robust estimation of the mounting angle for millimeter-wave automotive radars installed on moving vehicles is investigated. We propose a novel signal processing pipeline that combines radar and inertial measurement unit (IMU) data to achieve accurate and reliable performance in realistic driving scenarios. Unlike previous studies, the method employs neural networks to process sparse and noisy radar measurements, reject detections from moving objects, and estimate radar motion. In addition, a measurement model is introduced to correct IMU bias and scale factor errors. Using vehicle kinematics, the radar mounting angle is then computed from the estimated radar motion and the vehicle’s yaw rate. To benchmark performance, the proposed approach is comprehensively compared with two problem formulations and four estimation techniques reported in the literature. Validation is carried out on the challenging RadarScenes dataset, covering over 79 km of real-world driving. Results show that the proposed method achieves state-of-the-art accuracy and robustness, with reliable estimates obtained within approximately 25 seconds of driving. To the best of our knowledge, this is the first study to demonstrate that automotive radar mounting angles can be accurately estimated in complex, real traffic conditions, without requiring controlled environments, dedicated targets, or specially designed driving routes.
@@ I Introduction
Over the past decade, there has been increasing attention to and investment in advanced driver assistance systems (ADAS) and autonomous vehicles. To achieve safe and reliable autonomous driving, vehicles must be able to perceive complex environments with high accuracy. For this purpose, they are typically equipped with multiple perception sensors, including cameras, lidars, and automotive radars [1]. While cameras and lidars provide rich environmental information, their performance degrades significantly under adverse lighting and weather conditions [2]. In contrast, millimeter-wave (mmWave) automotive radar is more robust in such scenarios [3]. Combined with multiple-input multiple-output (MIMO) antenna technology, radar can achieve spatial resolution in compact hardware and provide object feature measurements such as range, azimuth, radial velocity, and elevation.
@@
Given these advantages, automotive radar has been widely adopted for tasks such as ego-motion estimation [4], environment mapping [5], moving object tracking [6], and road user classification [7]. However, these applications also rely on accurate knowledge of the radar’s extrinsic parameters, namely its mounting angle and position relative to the vehicle. Although such parameters are typically set at production, they can drift over time due to vibrations, collisions, or accidents. Even small deviations can severely degrade performance. For instance, as shown in Figure 1, a misalignment of only 0.05° in radar mounting angle can cause substantial localization errors. Errors can also be amplified by radar’s long detection range, degrading tasks such as mapping and sensor fusion [8]. Regular calibration of radar extrinsic parameters in operational driving conditions is therefore essential.
@@
Among extrinsic parameters, the mounting angle is the most critical and prone to change [8, 9]. Traditional approaches determine it using handheld compasses, angle sensors, or radar housings with accelerometers and actuators [10, 11, 12]. However, these methods are costly, labor-intensive, and often require skilled engineers, making them impractical for regular in-vehicle calibration. Recent research has therefore shifted toward algorithmic solutions [8], with the ultimate goal of fast, robust, and accurate mounting angle estimation that works automatically in real traffic conditions. Yet most existing methods have only been validated in controlled environments [13], with special targets [14], or on carefully designed driving routes [15].
To address these limitations, we propose a novel signal processing pipeline for radar mounting angle estimation in operational conditions. The method integrates radar and inertial measurement unit (IMU) data and exploits rigid-body kinematics to estimate the radar’s mounting angle. Key advantages include:
  1. 1.
    Robust sensing: Only radar and IMU are required, both of which remain reliable under adverse weather and lighting compared to optical sensors.
  2. 2.
    No controlled setup: Unlike previous studies, the method directly estimates mounting angle in real traffic, avoiding assumptions about controlled environments or specially designed routes.
  3. 3.
    Novel motion estimation: A neural network (NN) rejects detections from moving objects and estimates radar motion more robustly than popular RANSAC-based methods, even when moving objects dominate in the scene.
  4. 4.
    Accurate and fast convergence: The method provides stable estimates with mean absolute error <0.01deg<0.01\ deg and error variance $0.01\ deg^{2}$ within 25 seconds of driving, significantly faster than previous approaches.
@@
For comprehensive evaluation, this study examines two problem formulations and four estimation techniques within this framework. Experiments use the challenging RadarScenes dataset, covering more than 79 km of urban driving across varied traffic and environmental conditions. To the best of our knowledge, this is the first study to demonstrate accurate and robust estimation of radar mounting angles in unconstrained real-world traffic, without controlled environments, dedicated radar targets, or specific driving maneuvers.
@@ II Related Works
    Sensitivity to outliers. To handle complex driving scenarios, most odometry-based methods use random sample consensus (RANSAC) [29] or its variants [30] to mitigate the impact of measurement noise and moving objects (i.e., outliers) on radar motion estimation [28, 25, 24]. However, RANSAC is an iterative algorithm with several parameters to tune and assumes that most radar measurements originate from stationary objects. Its performance degrades significantly when outliers dominate, as in dense traffic with many moving targets. Consequently, many studies evaluate their methods only in controlled environments where most surrounding objects are static [31, 24, 25], leaving robustness in realistic driving scenarios insufficiently demonstrated.
@@
In summary, odometry-based methods offer clear advantages over other approaches. However, the diversity of road users and the dynamic behavior of vehicles hinder their robustness in real-world scenarios. Moreover, while different problem formulations and estimation techniques have been proposed [24, 28], a comprehensive comparison of their bottlenecks and trade-offs is still lacking.
@@ III-A Overview of Proposed Pipeline
The proposed processing pipeline belongs to the category of odometry-based methods and, more specifically, relies on an IMU. Its objective is to estimate the mounting (yaw) angle of an automotive radar system installed on a moving vehicle. Throughout this work, the vehicle motion is assumed to have no side-slip.
As illustrated in Figure  2, the pipeline takes radar point clouds and IMU yaw-rate measurements as inputs. The radar point cloud is processed by a NN-based radar motion estimator. In contrast to previous studies in this category, which mainly employ RANSAC-based radar motion estimators, the proposed approach is more robust to the scenarios with high outlier ratios in radar measurements. The NN-based radar motion estimator outputs radar motion vectors as well as estimated point weights [35]. These weights are subsequently used for variance estimation and for rejecting sparse frames.
@@
Figure 2: Overview of the proposed signal processing pipeline for radar mounting angle estimation. The inputs are radar point clouds and IMU yaw-rate readings, and the output is the estimated radar mounting angle. Here, $\mathbf{V}_{t}^{\mathrm{radar}}$ denotes the radar motion vector at timestamp $t$, $\omega_{t}$ is the vehicle yaw rate (rotational velocity), $(x_{s},y_{s})$ are the radar coordinates with respect to the vehicle rear center, $\beta_{t}$ is the direction of radar motion in the radar coordinate frame, and $\theta$ is the unknown radar mounting angle.
@@ III-B Radar Motion Estimation
The main objective of the radar motion estimator is to process raw radar point clouds and estimate the radar motion. In the literature, two main approaches exist for radar motion estimation: scan-matching methods [36] and instantaneous methods [4]. It has been shown in [35] that, under realistic driving scenarios, instantaneous methods [4, 35] are generally more robust than scan-matching methods and yield more accurate motion estimates. Furthermore, most approaches for radar mounting angle estimation rely on instantaneous methods [13, 21, 28, 25]. However, as discussed in Section II, these methods typically depend on random sampling (e.g., RANSAC [29] or its variants) for outlier rejection, and their performance degrades significantly when the radar point cloud contains a high proportion of outliers. In addition, RANSAC is an iterative algorithm, which can lead to poor runtime performance.
@@
For the multi-frame case, motion estimates accumulated over $T$ consecutive radar frames are processed by an additional temporal NN, denoted as the “NN-based Smoother” in Figure  3. This temporal NN captures the hidden relationship between non-zero vehicle acceleration and Doppler spectrum broadening, [^1] Vehicle acceleration causes Doppler migration and spectrum broadening, such that the measured radial velocity of a static object no longer matches the instantaneous vehicle motion. The magnitude of the mismatch depends on the acceleration. while also smoothing the initial estimates according to a second-order motion model. The output of the temporal NN consists of $T$ motion estimates $\mathbf{V}_{t\rightarrow t+T}^{\mathrm{radar}}$ and a weight matrix $\mathbf{W}_{J,t\rightarrow t+T}^{\mathrm{est}}$.
To further improve robustness, the estimated radar motion and point weights are used to compute the estimation variance and to reject sparse radar frames. First, the number of inlier points $L_{t}$ is calculated from the weights:
@@
where $IT$ denotes the inlier threshold. Using the angle $\alpha_{t}^{l}$ and Doppler $d_{t}^{l}$ measurements of the $L_{t}$ inliers, together with the estimated 2D radar motion $\mathbf{V}_{t}^{\mathrm{radar}}=\big[v_{x,t}^{\mathrm{radar}},\;v_{y,t}^{\mathrm{radar}}\big]^{\top}$, the residual error $\epsilon_{t}$ is computed as
@@ III-C Inertial Measurement Unit
The proposed method uses yaw-rate readings from an IMU. Compared with other odometry sensors such as GPS, IMUs are more robust in urban canyons and in the presence of tall buildings. Nevertheless, several factors limit the accuracy of IMU yaw-rate measurements. To account for these effects, this work adopts a standard yaw-rate measurement model [28] that includes a scale factor, a bias, and additive noise:
@@ III-D Problem Formulation
Linearizing $\beta_{t}(\theta,s^{\prime})$ at $(\theta_{0},s^{\prime}_{0})$ via first-order Taylor expansion yields
@@ III-E Mounting Angle Estimation
From the linearized expression in Equation 11, stacking $T$ radar frames yields an overdetermined linear system ($T>2$) in the unknowns $\theta$ (mounting angle) and $s^{\prime}$ (inverse IMU scale factor). Setting the linearization point $s^{\prime}_{0}=1$, the system can be written as
@@
Based on the covariance of the radar motion estimates, the wLSQ solution is
@@ IV-A Dataset and Evaluation Details
Figure 4: Density distribution of ego-vehicle speed and acceleration. For visualization, only 1% of the data (about 4.3k radar frames) was randomly sampled from the 64 selected recordings.
In addition to challenging data, this work also conducts a comprehensive performance comparison. First, the proposed method (both single-frame, SF, and multi-frame, MF) is compared with two representative methods from the literature (Table I). The comparison considers not only estimation accuracy, but also robustness and convergence speed. Moreover, this work evaluates the relative trajectory error (RTE) [^2] RTE measures the difference between two trajectories. It first aligns the starting points of the estimated and reference trajectories, then computes the $\ell_{2}$ distance between their end points (unit in meters). For longer trajectories, the path can be divided into shorter segments (e.g., a 2 km trajectory divided into 50 m segments). RTE is then computed on each segment, and the final reported error (e.g., RTE_50) is the mean over all segments. metric [35], which quantifies the impact of mounting angle misalignment on vehicle positioning and also highlights imperfections in the ground truth. Finally, this work explores alternative problem formulations and estimation techniques by modifying the proposed pipeline.
@@
IV-B Robustness
As discussed in Section II, many previous studies evaluate radar mounting angle estimation in controlled environments (e.g., parking lots), where most surrounding objects are stationary. In contrast, real driving scenarios involve numerous moving vehicles, diverse road conditions, and a wide range of speeds and accelerations. It is therefore crucial to assess whether the proposed method provides robust angle estimates under such challenging conditions. To this end, the estimation accuracy of all compared methods is evaluated across the 64 testing scenes. Figure 5 presents the estimated mounting angle for each scene. The two baseline methods from the literature show large fluctuations around the ground-truth angle, and their performance is clearly affected by the driving environment. For example, both methods produce large errors in Scene Index 36, which contains many moving vehicles in front of and near the ego-vehicle. Even after averaging over all 64 scenes, the mean estimates of the baselines remain far from the ground truth, with the smallest mean error still around $0.0444^{\circ}$.
In comparison, the proposed single-frame (SF) method is substantially more robust: it achieves an 80.0% reduction in error variance and a 69.8% improvement in mean accuracy compared with the best baseline. When multiple radar frames are available (Figure 5d), the proposed multi-frame (MF) method further exploits temporal motion correlations to mitigate the effects of outliers and non-zero vehicle acceleration, leading to even lower variance and higher robustness. It should be noted that all results in this section are based on data from a forward-looking radar (‘Radar 3’ in the dataset), which is particularly challenging because of the large number of moving objects in its FoV. Consequently, even with the proposed method, the estimates cannot be made fully consistent across all scenes, and the variance cannot be driven to zero. Preliminary experiments suggest that fusing data from multiple radars could alleviate this limitation, but a full exploration of multi-radar fusion is left for future work.

### v2

# Radar Mounting Angle Estimation in Operational Driving Conditions

@@ Abstract
The problem of estimating the mounting angle of millimeter-wave automotive radars installed on moving vehicles is investigated. We address this angle estimation problem during normal driving, without relying on controlled environments, dedicated radar targets, or specially designed driving routes. To achieve this, we propose a signal processing pipeline that combines radar and inertial measurement unit (IMU) data to enable accurate and reliable estimation under realistic driving conditions. Unlike previous studies, the method employs neural networks to process sparse and noisy radar measurements, reject detections from moving objects, and estimate radar motion. In addition, a measurement model is introduced to correct IMU bias and scale factor errors. Using vehicle kinematics, the radar mounting angle is then computed from the estimated radar motion and the vehicle’s yaw rate. To benchmark performance, the proposed approach is comprehensively compared with two alternative problem formulations and four estimation techniques reported in the literature. Validation is carried out on the challenging RadarScenes dataset, covering over 79 km of real-world driving with different velocities and trajectories. Results show that stable and accurate mounting angle estimates are obtained within approximately 25 seconds of driving. To the best of our knowledge, this is the first study to demonstrate that automotive radar mounting angles can be estimated during complex, real traffic conditions using only onboard sensor data.
@@ I Introduction
Over the past decade, there has been increasing attention to advanced driver assistance systems (ADAS) and autonomous vehicles. To achieve safe and reliable autonomous driving, vehicles must be able to perceive complex environments with high accuracy. For this purpose, they are typically equipped with multiple perception sensors, including cameras, lidars, and automotive radars [12]. While cameras and lidars provide rich environmental information, their performance degrades significantly under adverse lighting and weather conditions [11]. In contrast, millimeter-wave (mmWave) automotive radar maintains reliable sensing performance under such conditions and has therefore become a key sensing modality for automotive perception [18]. Combined with multiple-input multiple-output (MIMO) antenna technology, radar can achieve spatial resolution in compact hardware and provide object feature measurements such as range, azimuth, radial velocity, and elevation.
@@
Given these advantages, automotive radar has been widely adopted for tasks such as ego-motion estimation [22], environment mapping [35], moving object tracking [10], and road user classification [31]. A common prerequisite for all these applications is accurate knowledge of the radar’s extrinsic parameters, in particular its mounting angle relative to the vehicle. Although such parameters are typically set at the production stage, they can drift over time due to vibrations, collisions, or accidents. Even small deviations can severely degrade performance. For instance, as shown in Figure 1, a misalignment of only 0.05° in radar mounting angle can cause substantial localization errors. Errors can also be amplified by radar’s long detection range, degrading crucial tasks such as mapping and sensor fusion [5]. Regular estimation of the radar mounting angle during operational driving conditions is therefore very important [38].
@@
Traditional approaches determine this angle using handheld compasses, angle sensors, or radar housings with accelerometers and actuators [16, 4, 29]. However, these methods are costly, labor-intensive, and often require skilled engineers, making them impractical for regular in-vehicle calibration. Recent research has therefore shifted toward algorithmic solutions [5], which rely on measurements from the radar under calibration together with an additional reference sensor. While these approaches aim to enable fast and accurate mounting angle estimation that operates automatically under normal driving conditions, most existing methods have only been validated in controlled environments [14], with special targets [27], or on carefully designed driving routes [8].
In practice, the key challenge in radar mounting angle estimation is not the formulation of the estimation formula itself, but the quality of the available radar measurements under operational driving conditions. In dense traffic, a large fraction of radar detections originate from moving objects unsuitable to be used as references for calibration purposes; vehicle acceleration affects Doppler estimation; and radar point clouds can be sparse and flickering. These factors collectively cause existing calibration methods to fail unless restrictive assumptions, such as static scenes or controlled routes, are imposed.
@@
This work addresses this bottleneck by formulating a mounting angle estimation approach that remains applicable under normal, uncontrolled driving conditions. Based on the proposed formulation, the approach combines complementary onboard sensor measurements within a kinematic framework to enable reliable radar mounting angle estimation under challenging real-world measurement conditions. For comprehensive evaluation, this study examines two problem formulations and four estimation techniques within this framework. Experiments use the challenging RadarScenes dataset [33], covering more than 79 km of urban driving across varied traffic and environmental conditions, as well as varying velocity and trajectories of the ego-vehicle. To the best of our knowledge, this study demonstrates that radar mounting angles can be accurately estimated in unconstrained real-world traffic, without relying on controlled environments, dedicated radar targets, or specific driving maneuvers.
@@ II Related Works
    Sensitivity to outliers. To handle complex driving scenarios, most odometry-based methods use random sample consensus (RANSAC) [13] or its variants [37] to mitigate the impact of measurement noise and moving objects (i.e., outliers) on radar motion estimation [21, 9, 1]. However, RANSAC is an iterative algorithm with several parameters to tune and assumes that most radar measurements originate from stationary objects. Its performance degrades significantly when outliers dominate, as in dense traffic with many moving targets. Consequently, many studies evaluate their methods only in controlled environments where most surrounding objects are static [20, 1, 9], leaving performance under realistic driving scenarios less extensively validated.
@@
In summary, odometry-based methods offer clear advantages over other approaches. However, the diversity of road users and the dynamic behavior of vehicles pose significant challenges for their application in real-world driving scenarios. Moreover, while different problem formulations and estimation techniques have been proposed [1, 21], a comprehensive comparison of their bottlenecks and trade-offs is still lacking.
@@ III-A Overview of Proposed Pipeline
The proposed processing pipeline belongs to the category of odometry-based methods and uses an inertial measurement unit (IMU) as the reference sensor. The objective is to estimate the mounting (yaw) angle of an automotive radar system installed on a moving vehicle. The formulation is based on rigid-body kinematics and assumes normal vehicle motion without lateral side-slip, which is a standard assumption in radar odometry literature [22, 34]. As illustrated in Figure 2, the pipeline takes radar point clouds and IMU yaw-rate measurements as inputs. The radar point cloud is processed by a neural network (NN) based radar motion estimator, which outputs radar ego-motion estimates together with estimated point weights [40]. The estimated weights are then used for variance estimation and for rejecting sparse radar frames.
Figure 2: Overview of the proposed signal processing pipeline for the problem of radar mounting angle estimation. Radar point clouds and IMU yaw-rate measurements are used as inputs, and the radar mounting angle is estimated as output. Radar motion is first estimated from the point clouds, while IMU yaw-rate measurements are modeled to account for bias and scale factor effects. The mounting angle is then obtained by enforcing a lateral velocity equality constraint within a kinematic formulation and solving a weighted least-squares problem. Here, $\mathbf{V}_{t}^{\mathrm{radar}}$ denotes the radar motion vector at timestamp $t$, $\omega_{t}$ is the vehicle yaw rate (rotational velocity), $(x_{s},y_{s})$ are the radar coordinates with respect to the vehicle rear center, $\beta_{t}$ is the direction of radar motion in the radar coordinate frame, and $\theta$ is the unknown radar mounting angle.
@@
It is important to highlight that the overall structure of the proposed pipeline is driven by the practical constrains encountered in real-world driving. Adverse weather and lighting conditions limit the reliability of optical sensors, motivating the use of an IMU as the reference sensor due to its robustness to environmental conditions, high refresh rate, and low data throughput. At the same time, radar point clouds captured in dense traffic are dominated by measurements from moving objects and are often sparse or affected by vehicle acceleration, which makes classical model-based approaches unreliable. A learning-based method is therefore employed to process radar data, handle these unfavorable conditions, and estimate radar motion. Finally, a kinematic formulation is adopted to relate radar motion and vehicle rotation without introducing vehicle-specific dynamic parameters. Together, these components form a compact pipeline required to estimate the radar mounting angle under operational driving conditions.
@@ III-B Radar Motion Estimation
The main objective of the radar motion estimator is to process raw radar point clouds and estimate the radar motion. In the literature, two main approaches exist for radar motion estimation: scan-matching methods [30] and instantaneous methods [22]. It has been shown in [40] that, under realistic driving scenarios, instantaneous methods [22, 40] are generally less sensitive to point cloud sparsity and dynamic objects than scan-matching methods. As a result, most existing approaches for radar mounting angle estimation rely on instantaneous methods [14, 7, 21, 9]. However, as discussed in Section II, these methods typically depend on random sampling (e.g., RANSAC [13] or its variants) for outlier rejection, and their performance degrades significantly when the radar point cloud contains a high proportion of outliers. In addition, RANSAC is an iterative algorithm, which can lead to poor runtime performance.
@@
For the multi-frame case, motion estimates accumulated over $T$ consecutive radar frames are processed by an additional temporal NN, denoted as the “NN-based Smoother” in Figure 3. This temporal NN captures the hidden relationship between non-zero vehicle acceleration and Doppler spectrum broadening, [^1] Vehicle acceleration causes Doppler migration and spectrum broadening, such that the measured radial velocity of a static object no longer matches the instantaneous vehicle motion. The magnitude of the mismatch depends on the acceleration. while also smoothing the initial estimates according to a second-order motion model. The output of the temporal NN consists of $T$ motion estimates $\mathbf{V}_{t\rightarrow t+T}^{\mathrm{radar}}$ and a weight matrix $\mathbf{W}_{J,t\rightarrow t+T}^{\mathrm{est}}$.
To further improve estimation reliability, the estimated radar motion and point weights are used to compute the estimation variance and to reject sparse radar frames. First, the number of inlier points $L_{t}$ is calculated from the weights:
@@
where $IT$ denotes the inlier threshold. Using the angle $\alpha_{t}^{l}$ and Doppler $d_{t}^{l}$ measurements of the $L_{t}$ inliers, together with the estimated 2D radar motion $\mathbf{V}_{t}^{\mathrm{radar}}=\big[v_{x,t}^{\mathrm{radar}},\;v_{y,t}^{\mathrm{radar}}\big]^{\top}$, the residual error $\epsilon_{t}$ is computed as:
@@ III-C Inertial Measurement Unit
The proposed method uses yaw-rate readings from an IMU. Compared with other odometry sensors such as GPS, IMUs provides more reliable measurements in urban areas and in the presence of tall buildings. Nevertheless, several factors limit the accuracy of IMU yaw-rate measurements. To account for these effects, this work adopts a standard yaw-rate measurement model [21] that includes a scale factor, a bias, and additive noise:
@@ III-D Problem Formulation
Linearizing $\beta_{t}(\theta,s^{\prime})$ at $(\theta_{0},s^{\prime}_{0})$ via first-order Taylor expansion yields:
@@ III-E Mounting Angle Estimation
From the linearized expression in Equation 11, stacking $T$ radar frames yields an overdetermined linear system ($T>2$) in the unknowns $\theta$ (mounting angle) and $s^{\prime}$ (inverse IMU scale factor). Setting the linearization point $s^{\prime}_{0}=1$, the system can be written as:
@@
Based on the covariance of the radar motion estimates, the wLSQ solution is:
@@
Finally, it is worth noting that the proposed method relies on ordinary vehicle motion to enable mounting angle estimation. In particular, non-zero vehicle rotation is required so that the lateral velocity induced by vehicle yaw motion can be observed at the radar location. In addition, forward motion of the ego-vehicle is also necessary. The above conditions can be satisfied during normal driving operations, without the need of specially designed maneuvers or controlled motion patterns.
@@ IV-A Dataset and Evaluation Details
Figure 4: Density distribution of ego-vehicle speed and acceleration. For visualization, only 1% of the data (about 4.3k radar frames) was randomly sampled from the 64 selected recordings. This figure shows that the dataset used for evaluation covers a wide range of vehicle motion states during normal driving operations.
In addition to challenging data, this work also conducts a comprehensive performance comparison. First, the proposed method (both single-frame, SF, and multi-frame, MF) is compared with two representative methods from the literature (Table I). This comparison considers not only estimation accuracy, but also estimation stability and convergence speed. Moreover, this work evaluates the relative trajectory error (RTE) [^2] RTE measures the difference between two trajectories. It first aligns the starting points of the estimated and reference trajectories, then computes the $\ell_{2}$ distance between their end points (unit in meters). For longer trajectories, the path can be divided into shorter segments (e.g., a 2 km trajectory divided into 50 m segments). RTE is then computed on each segment, and the final reported error (e.g., RTE_50) is the mean over all segments. metric [40], which quantifies the impact of mounting angle misalignment on vehicle positioning and also highlights imperfections in the ground truth. Finally, this work explores alternative problem formulations and estimation techniques by modifying the proposed pipeline.
@@
IV-B Performance Across Diverse Driving Scenes
As discussed in Section II, many previous studies evaluate radar mounting angle estimation in controlled environments (e.g., parking lots), where most surrounding objects are stationary. In contrast, real driving scenarios involve numerous moving vehicles, diverse road conditions, and a wide range of vehicle speeds and accelerations. It is therefore important to evaluate how mounting angle estimates behave across different driving scenes with varying levels of scene complexity. To this end, the estimation accuracy of all compared methods is evaluated across the 64 testing scenes. Figure 5 presents the estimated mounting angle for each scene. The two baseline methods from the literature show large fluctuations around the ground-truth angle, and their performance is clearly affected by the driving environment. For example, both methods produce large errors in Scene Index 36, which contains many moving vehicles in front of and near the ego-vehicle. Even after averaging over all 64 scenes, the mean estimates of the baselines remain far from the ground truth, with the smallest mean error still around $0.0444^{\circ}$.
In comparison, the proposed single-frame (SF) method shows substantially reduced variability across scenes, achieving an 80.0% reduction in error variance and a 69.8% improvement in mean accuracy compared with the best baseline. When multiple radar frames are available (Figure 5d), the proposed multi-frame (MF) method further exploits temporal motion correlations to mitigate the effects of outliers and non-zero vehicle acceleration, resulting in more consistent estimates across scenes. It should be noted that all results in this section are based on data from a forward-looking radar (‘Radar 3’ in the dataset), which is particularly challenging because of the large number of moving objects in its FoV. Consequently, even with the proposed method, the estimates cannot be made fully consistent across all scenes, and the variance cannot be driven to zero. Preliminary experiments suggest that fusing data from multiple radars could alleviate this limitation, but a full exploration of multi-radar fusion is left for future work.